# The Developing Human Connectome Project: A Fast Deep Learning-based Pipeline for Neonatal Cortical Surface Reconstruction

### v1

@@ 2.6 Cortical Surface Inflation
Following the HCP [25] and original dHCP pipelines [44], our dHCP DL pipeline incorporates two types of cortical surface inflation. Firstly, the original pipeline employs Connectome Workbench [25] to iteratively inflate and smooth the midthickness surface to generate inflated and very inflated surfaces. These surfaces are primarily used for visualization. Our pipeline re-implements these surface inflation approaches using PyTorch to facilitate GPU acceleration.
Secondly, the original dHCP pipeline adopts FreeSurfer mris_inflate command [22, 21], which is reproduced in the Medical Image Registration ToolKit (MIRTK) [56, 44], to inflate the white matter surface while minimizing the metric distortion between white matter and inflated surfaces. During the inflation, the cortical sulcal depth, or average convexity for each mesh vertex is computed by integrating the displacement of the vertex along its normal direction. Our DL-based pipeline also accelerates such an inflation process by PyTorch with parallel computation on a GPU.
@@ 2.7 Cortical Feature Estimation
For the myelin map estimation, if the T1 image is provided and aligned to the T2 image, we divide the T1 image by the original T2 image before intensity normalization to compute the T1/T2 ratio. Then, the myelin map is estimated by projecting the T1/T2 ratio onto the midthickness surface using volume-to-surface mapping [26, 25, 44], which is implemented with the Connectome Workbench. For each vertex of the midthickness surface mesh, we find its corresponding location in the T1/T2 ratio volume, and compute the Gaussian average of its neighboring voxels as the myelin map value, where the size of the neighbor depends on the cortical thickness. Only voxels within the cortical ribbon ROI are considered. To accelerate the volume-to-surface mapping, we train a 3D U-Net to learn the binary cortical ribbon mask, i.e., the cortical GM segmentation. The learning procedure is similar to brain extraction as shown in Figure 4. The GT segmentation is generated by the Draw-EM approach integrated in the original dHCP pipeline.
@@ 2.8 Spherical Projection
The purpose of spherical projection or spherical mapping is to project the white matter surface onto a sphere while minimizing metric distortions. Motivated by [70], we propose an unsupervised learning framework for spherical projection and integrate it into our DL-based dHCP pipeline. As illustrated in Figure 7, the inputs of our spherical projection framework include a white matter surface and a fixed initial sphere for all subjects. Note that since all white matter surfaces are deformed from the same initial template surface, the white matter surface has the same mesh connectivity as the initial sphere, which is generated from the initial surface by FreeSurfer (see Figure 3). Therefore, instead of inflating the white matter surface and projecting the inflated surface to create a sphere for each subject [22, 18, 70], our approach learns to deform the initial sphere directly while minimizing metric distortions between the deformed sphere and the white matter surface. In this work, we primarily consider edge and area distortions.
We resample the input white matter surface mesh such that it has the same connectivity as a standard icosphere with 163,842 vertices. This is achieved by barycentric interpolation [52, 51] between the initial sphere and the icosphere, of which the barycentric coordinates are pre-computed to accelerate the resampling. Then, we compute the metrics of the resampled white matter surface including edge length and face area. Each vertex contains the metrics of its adjacent edges and faces. We concatenate the coordinates of the resampled white matter surface and the icosphere as well as computed metrics as input features of a Spherical U-Net [71], which has been widely used in cortical surface analysis tasks such as cortical surface registration [68], parcellation [72] and development prediction [69]. We train the Spherical U-Net to learn a diffeomorphic deformation field in the spherical domain by integrating a SVF using scaling and squaring [2]. Rather than adding regularization terms to the loss function [70], we apply Laplacian mesh smoothing to the deformation field to encourage smoothness. The predicted sphere is derived by deforming the initial sphere, where the displacement of each vertex is determined by sampling from the deformation field using barycentric interpolation.
@@
3 Comparison to Existing Neuroimage Pipelines
3.1 Comparison to Original dHCP Pipeline
@@
3.1.1 Approach and Implementation
3.1.2 Runtime
To measure the runtime, we run both original and DL-based dHCP pipeline to process the brain MRI of 10 dHCP neonatal subjects, which are randomly selected from the test set. The runtime is measured on an 8-core Intel Core i7-11700K CPUs and a NVIDIA GeForce RTX 3080 GPU with 12GB memory. The original dHCP pipeline [44] runs exclusively on CPUs and requires 6h38min to process a single subject. The volume-based and surface-based processing requires 2h9min and 4h29min respectively. For our DL-based pipeline, we report the runtime with and without GPU acceleration in Table 2. With GPU acceleration, it only takes 24 seconds to run the entire DL-based pipeline, which is 995$\times$ faster than the original dHCP pipeline. The cortical surface reconstruction and spherical projection only needs 3 seconds in total owing to the powerful end-to-end deep learning approaches. Even without GPU acceleration, our DL-based pipeline only requires 170 seconds to complete on the CPU, which is still 140$\times$ faster compared to the original dHCP pipeline.
@@ 3.1.2 Runtime
Table 3: Comparisons of functionality and runtime among DL-based neuroimage processing pipelines: FastSurfer, iBEAT V2.0, and our dHCP DL pipeline. The green check marks indicate DL-based processing steps. We provide the approximated runtime as reported in original papers.
Fig. 8: The results of quality control (QC). The assessors manually examine and compare the quality of cortical surfaces between the original and our DL-based dHCP pipelines for each test sample. Top: the average QC results. For 82.5% of all test samples, our DL-based dHCP pipeline achieves superior (54.2%) or equal (28.3%) surface quality compared to the original dHCP pipeline. Bottom: the individual QC results from all assessors.
Fig. 9: The cortical morphological features including cortical volume ($cm^{3}$), surface area ($m^{2}$), average cortical thickness ($mm$), and average absolute value of sulcal depth ($mm$) across 265 test neonatal subjects with different PMA at scan. The Pearson correlation coefficients (PCC) $r$ are reported to measure the correlation between the morphological features and the PMA.
Fig. 10: The qualitative comparison between original and our DL-based dHCP pipeline on different test samples of the dHCP dataset. The left columns visualize the corruptions in cortical surfaces produced by the original dHCP pipeline, while the right columns show that our DL-based dHCP pipeline effectively reduces these errors. Row 1-2: corruptions on the top of the cortical surfaces. Row 2-4: corruptions on the back of the cortical surfaces. Row 5: the ”cave/hole” introduced by the errors of WM segmentations.
@@
Fig. 12: The qualitative comparison between original and our DL-based dHCP pipeline on two neonatal subjects of the ePrime dataset. The cortical surfaces generated by the original dHCP pipeline are severely corrupted and broken on gyri, whereas our DL-based pipeline has better generalization ability. From top to bottom: T2w MRI, left white matter surface, left pial surface, right white matter surface, and right pial surface.
@@ 4.1.1 Visual Inspection
We conduct manual quality control (QC) via visual inspection to assess the quality of the cortical surfaces generated by our DL-based dHCP pipeline. More specifically, our QC aims to examine if our DL-based dHCP pipeline produces better cortical surfaces compared to the baseline of the original dHCP pipeline [44].
@@
We randomly select 30 T2w brain MRI samples from 265 test cases and extract two groups of cortical surfaces for each sample by the original dHCP pipeline and our DL-based dHCP pipeline respectively. QC is performed by four expert assessors with clinical background or extensive experience in neuroimage analysis. To ensure fairness, all assessors are requested to visually evaluate two groups of surfaces independently, and select the group with better quality for each test sample without knowing the corresponding reconstruction approaches (original or DL-based pipeline). Equal quality is also an option when both groups of surfaces are anatomically accurate or corrupted to the same degree. The average and individual QC results are reported in Figure 8. The results show that for 82.5% of all test samples, our DL-based dHCP pipeline achieves superior (54.2%) or equal (28.3%) surface quality compared to the original dHCP pipeline. The individual QC results in Figure 8 also indicate that our DL-based dHCP pipeline achieves better performance in the evaluation of all assessors. These demonstrate that our DL-based dHCP pipeline reconstructs higher quality cortical surfaces, while being 995$\times$ faster than the original dHCP pipeline.
@@
We further provide visual comparison between cortical surfaces reconstructed by original and DL-based dHCP pipeline. Figure 10 indicates that the original dHCP pipeline is prone to produce corruptions on the top and back (parietal and occipital lobes) of the cortical surfaces, due to the low image contrast in the corresponding regions of T2w brain MRI, whereas our DL-based pipeline achieves relatively high anatomical accuracy in these regions. In addition, the surfaces generated by the original dHCP pipeline are affected by the imprecise Draw-EM segmentations [43, 44], and thus occasionally produce unexpected ”caves” or ”holes” as shown in Figure 10. Our DL-based pipeline leverages diffeomorphic surface deformations and learns cortical surfaces end-to-end from brain MRI, which can effectively avoid such issues.
@@ 4.1.2 Brain Development
In addition to manual quality control, we provide qualitative and quantitative evaluation to validate that the cortical surfaces extracted by our dHCP DL pipeline can capture the brain development of neonates. We visualize the cortical surfaces as well as the corresponding cortical features of different neonatal subjects scanned at the PMA of 29,34,39,44 weeks in Figure 11. The cortical folding is observed in Figure 11 with the increasing of the PMA. The cortical thickness and sulcal depth are increasing as well with brain growth.
We further measure and plot cortical morphological features including cortical volume, surface area, cortical thickness and sulcal depth across all 265 test neonatal subjects in Figure 9. The cortical volume is equal to the volumes of pial surfaces minus those of the white matter surfaces for both brain hemispheres. The surface area refers to the mesh area of the pial surfaces. The cortical thickness and sulcal depth for each subject is measured by averaging the cortical thickness and the absolute value of the sulcal depth over all vertices on the surface meshes. Figure 9 shows the positive trends of the morphological features with increasing PMA, which reflect the growth and folding of the neonatal cerebral cortex. Moreover, we compute the Pearson correlation coefficients (PCC) $r$ to measure the correlation between the morphological features and the PMA. The PCCs are provided in Figure 9, which indicate that all cortical morphological features have strong positive correlation with the age of the neonates. This verifies that the cortical surfaces extracted by our DL-based pipeline are able to capture the brain development of neonates.
@@
Table 4: The number of self-intersection faces (SIFs) on cortical surfaces.

### v2

@@ 2.6 Cortical Surface Inflation
Following the HCP [26] and original dHCP pipelines [50], our dHCP DL pipeline incorporates two types of cortical surface inflation. Firstly, the original pipeline employs Connectome Workbench [26] to iteratively inflate and smooth the midthickness surface to generate inflated and very inflated surfaces. These surfaces are primarily used for visualization. Our pipeline re-implements these surface inflation approaches using PyTorch to facilitate GPU acceleration. Secondly, the original dHCP pipeline adopts FreeSurfer mris_inflate command [23, 22], which is reproduced in the Medical Image Registration ToolKit (MIRTK) [62, 50], to inflate the white matter surface while minimizing the metric distortion between white matter and inflated surfaces. Our DL-based pipeline also accelerates such an inflation process by PyTorch with parallel computation on a GPU.
During the inflation, the cortical sulcal depth, or average convexity of each mesh vertex is computed by integrating the displacement of the vertex along its normal direction. As all cortical surfaces have fixed number of vertices, our DL-based dHCP pipeline provides age-consistent estimation of sulcal depth, while original dHCP pipeline extracts cortical surfaces with less number of vertices for younger infants, resulting in over-inflation of surfaces and overestimation of sulcal depth.
@@ 2.7 Cortical Feature Estimation
For the myelin map estimation, if the T1 image is provided and aligned to the T2 image, we divide the T1 image by the original T2 image before intensity normalization to compute the T1/T2 ratio. Then, the myelin map is estimated by projecting the T1/T2 ratio onto the midthickness surface using volume-to-surface mapping [27, 26, 50], which is implemented with the Connectome Workbench. For each vertex of the midthickness surface mesh, we find its corresponding location in the T1/T2 ratio volume, and compute the Gaussian average of its neighboring voxels as the myelin map value, where the size of the neighbor depends on the cortical thickness. Only voxels within the cortical ribbon ROI are considered. However, it is computationally intensive to examine if a voxel or a point is located between the inner and outer cortical surfaces, in particular for high resolution meshes with 150k vertices. To accelerate the volume-to-surface mapping, we train a 3D U-Net to learn the binary cortical ribbon mask, i.e., the cortical GM segmentation. The learning procedure is similar to brain extraction as shown in Figure 4. The pseudo GT segmentation is generated by the Draw-EM approach integrated in the original dHCP pipeline.
@@ 2.8 Spherical Projection
The purpose of spherical projection or spherical mapping is to project the white matter surface onto a sphere while minimizing metric distortions. Motivated by [77], we propose an unsupervised learning framework for spherical projection and integrate it into our DL-based dHCP pipeline. As illustrated in Figure 7, the inputs of our spherical projection framework include a white matter surface and a fixed initial sphere for all subjects. Note that since all white matter surfaces are deformed from the same initial template surface, the white matter surfaces have the same mesh connectivity as the initial sphere, which is generated from the initial surface by FreeSurfer (see Figure 3). Therefore, instead of inflating the white matter surface and projecting the inflated surface to create a sphere for each subject [23, 19, 77], our approach learns to deform the initial sphere directly while minimizing metric distortions between the deformed sphere and the white matter surface. In this work, we primarily consider edge and area distortions.
We resample the input white matter surface mesh such that it has the same connectivity as a standard icosphere with 163,842 vertices. This is achieved by barycentric interpolation [58, 57] between the initial sphere and the icosphere, of which the barycentric coordinates are pre-computed to accelerate the resampling. Then, we compute the metrics of the resampled white matter surface including edge length and face area. Each vertex contains the metrics of its adjacent edges and faces. We concatenate the coordinates of the resampled white matter surface and the icosphere as well as computed metrics as input features of a Spherical U-Net [78], which has been widely used in cortical surface analysis tasks such as cortical surface registration [75], parcellation [79] and development prediction [76]. We train the Spherical U-Net to learn a diffeomorphic deformation field in the spherical domain by integrating a SVF using scaling and squaring [33, 2]. Rather than adding regularization terms to the loss function [77], we apply Laplacian mesh smoothing to the deformation field to encourage smoothness. The predicted sphere is derived by deforming the initial sphere, where the displacement of each vertex is determined by sampling from the deformation field using barycentric interpolation.
@@
CPU Specs (Intel) Xeon Core Core
Gold 6154 i7-8700k i7-11700K
GPU Specs (NVIDIA) Titan Xp 1080 Ti 3080
Table 3: Comparisons of functionality and runtime among DL-based neuroimage processing pipelines: FastSurfer, iBEAT V2.0, and our dHCP DL pipeline. The green check marks indicate DL-based processing steps. We provide the approximated runtime as reported in original papers [32, 71]. The CPU and GPU specifications used in the original papers are reported as well for a fair comparison.
3 Comparison to Existing Neuroimage Pipelines
3.1 Comparison to Original dHCP Pipeline
3.1.1 Approach and Implementation
3.1.2 Runtime
To measure the runtime, we run both original and DL-based dHCP pipeline to process the brain MRI of 10 dHCP neonatal subjects, which are randomly selected from the test set. The runtime is measured on an 8-core Intel Core i7-11700K CPUs and a NVIDIA GeForce RTX 3080 GPU with 12GB memory. The original dHCP pipeline [50] runs exclusively on CPUs and requires 6h38min to process a single subject. The volume-based and surface-based processing requires 2h9min and 4h29min respectively. For our DL-based pipeline, we report the runtime with and without GPU acceleration in Table 2. With GPU acceleration, it only takes 24 seconds to run the entire DL-based pipeline, which is 995$\times$ faster than the original dHCP pipeline. The cortical surface reconstruction and spherical projection only need 3 seconds in total owing to the powerful end-to-end deep learning approaches. Even without GPU acceleration, our DL-based pipeline only requires 170 seconds to complete on the CPU, which is still 140$\times$ faster compared to the original dHCP pipeline.
Fig. 8: The results of manual cortical surface quality control (QC). The assessors visually examine and compare the anatomical correctness of cortical surfaces between the original and our DL-based dHCP pipelines for each test brain MRI. Top: the average surface QC results. For 82.5% of all test samples, our DL-based dHCP pipeline achieves superior (54.2%) or equal (28.3%) surface quality compared to the original dHCP pipeline. Bottom: the individual surface QC results from all assessors.
Fig. 9: The qualitative comparison between original and our DL-based dHCP pipeline on 5 different test samples of the dHCP dataset. The left columns visualize the corruptions in cortical surfaces produced by the original dHCP pipeline, while the right columns show that our DL-based dHCP pipeline effectively reduces these errors. The yellow boundary is the white matter surface, and the green boundary is the pial surface. Row 1-2: corruptions on the top of the cortical surfaces. Row 2-4: corruptions on the back of the cortical surfaces. Row 5: the ”cave/hole” introduced by the errors of WM segmentations.
Fig. 10: The number of test samples grouped by the visual QC scores.
Fig. 11: The visualization of succeeded and failed cases compared between original dHCP pipeline and our DL-based pipeline. Row 1-2: cortical surfaces of two test samples when 4 assessors agree that our DL-based pipeline produces better surface quality. Row 3-4: cortical surfaces of two test samples when 4 assessors agree that the surface quality of the original dHCP pipeline is no worse than ours.
Fig. 12: The visualization of surface distances from the cortical surfaces reconstructed by our DL-based dHCP pipeline to those generated by the original dHCP pipeline. The averaged cortical surfaces are displayed and overlaid with surface distance maps ($0.1-0.2mm$) averaged over all test samples. Top: the averaged white matter surfaces and distance maps. Bottom: the averaged pial surfaces and distance maps.
@@ 3.1.2 Runtime
Fig. 14: Comparison on cortical morphological features including cortical volume ($cm^{3}$), surface area ($m^{2}$), average cortical thickness ($mm$), and average absolute value of sulcal depth ($mm$) across 265 test neonatal subjects with different PMA at scan. The Pearson correlation coefficients (PCC) $r$ are reported to measure the correlation between the morphological features and the PMA. Left: cortical morphological features estimated by the original dHCP pipeline. Right: cortical morphological features estimated by our DL-based dHCP pipeline.
Fig. 15: The qualitative comparison between original and our DL-based dHCP pipeline on two neonatal subjects of the ePrime dataset. The cortical surfaces generated by the original dHCP pipeline are severely corrupted and broken on gyri, whereas our DL-based pipeline has better generalizability. From top to bottom: T2w MRI, left white matter surface, left pial surface, right white matter surface, and right pial surface.
@@ 4.1.1 Visual Inspection
In this section, we conduct manual quality control (QC) via visual inspection to assess the quality of the cortical surfaces generated by our DL-based dHCP pipeline. The cortical surface quality, i.e., the anatomical correctness of cortical surfaces, is measured by the number and severity of anatomical errors and corruptions in cortical surfaces. Due the lack of GT surfaces, we use the cortical surface generated by the original dHCP pipeline [50] as the baseline. More specifically, our visual QC aims to examine if our DL-based dHCP pipeline extracts cortical surfaces with better surface quality compared to the original dHCP pipeline.
@@
We randomly select 30 T2w brain MRI samples for QC from 265 test cases. For each test brain MRI, we extract two groups of cortical surfaces by original dHCP pipeline and our DL-based dHCP pipeline respectively. The visual QC is performed by four expert assessors with clinical background or extensive experience in neuroimage analysis to mitigate individual variability. To ensure fairness, all assessors visually evaluate two groups of cortical surfaces independently without knowing the corresponding reconstruction approaches (either original or DL-based pipeline). The assessors are requested to carefully examine and detect corrupted areas in two groups of cortical surfaces, and then select the group with better surface quality for each test sample based on the number and severity of identified defects and artifacts. Equal quality is also an option when both groups of cortical surfaces are anatomically correct or corrupted to the same degree.
The average and individual QC results are reported in Figure 8. The results show that for 82.5% of all test samples, our DL-based dHCP pipeline achieves superior (54.2%) or equal (28.3%) surface quality compared to the original dHCP pipeline in visual assessment. The individual QC results in Figure 8 also indicate that our DL-based dHCP pipeline achieves better performance in the qualitative evaluation of all assessors. These demonstrate that our DL-based dHCP pipeline reconstructs higher quality cortical surfaces, while being 995$\times$ faster than the original dHCP pipeline.
@@
Additionally, we provide visual comparison between cortical surfaces reconstructed by original and DL-based dHCP pipeline. Figure 9 displays the cortical surfaces of 5 different samples. It indicates that the original dHCP pipeline is prone to produce corruptions on the top and back (parietal and occipital lobes) of the cortical surfaces, due to the low image contrast in the corresponding regions of T2w brain MRI, whereas our DL-based pipeline achieves better anatomical correctness in these regions compared to the original dHCP pipeline. In addition, the surfaces generated by original dHCP pipeline are affected by the imprecise Draw-EM segmentations [49, 50], and thus occasionally produce unexpected ”caves” or ”holes” as shown in Figure 9. Our DL-based dHCP pipeline leverages diffeomorphic surface deformations and learns cortical surfaces end-to-end from brain MRI, which effectively avoids such issues.
Based on the manual QC results, we further measure the visual QC score for each of 30 test samples. The visual QC score $s$ of each test subject is computed by $s=\sum_{i=1}^{4}s_{i}$, where $s_{i}=1$ if the $i$-th assessor observes that our DL-based pipeline produces better surface quality compared to the baseline of the original dHCP pipeline, $s_{i}=-1$ if the original pipeline performs better, and $s_{i}=0$ means equal surface quality. For example, $s=4$ indicates all 4 assessors agree that our DL-based pipeline has better performance on surface quality. We count the number of test subjects grouped by the visual QC scores $-4\leq s\leq 4$. The results presented in Figure 10 show that 80% of the 30 test cases have positive visual QC scores $s>0$, which means our DL-based pipeline achieves better anatomical correctness than the original dHCP pipeline.
To explore both succeeded and failed cases of the proposed DL-based dHCP pipeline, in Figure 11, we visualize the test samples with QC scores $s=4$ and $s=-3$, i.e., all 4 assessors agree that the surface quality of our DL-based dHCP pipeline is better and no better than the original pipeline. Row 1-2 in Figure 11 shows that the succeeded cases ($s=4$) can alleviate the systematic errors on the top and back of the cortical surfaces produced by the original dHCP pipeline. However, the failed examples ($s=-3$) of our DL-based pipeline in Row 3-4 still introduce corruptions in the occipital lobe, affected by the limited surface quality of the pseudo GT cortical surfaces extracted by the original dHCP pipeline.
4.1.2 Quantitative Evaluation
In addition to visual inspection, we conduct quantitative evaluation to validate that our DL-based dHCP pipeline can alleviate the anatomical errors produced by original dHCP pipeline [50] in areas with low image contrast, e.g., the top and back (occipital lobe) of the brain. Following previous work [74, 41, 44], we compute the surface distances from the vertices of the cortical surfaces predicted by our DL-based pipeline to the pseudo GT cortical surfaces generated by the original pipeline. Since the cortical surfaces of our DL-based pipeline share the same mesh connectivity, we visualize the averaged cortical surfaces overlaid with the surface distance maps averaged over all 265 test cases in Figure 12.
The distance maps displayed in Figure 12 show large surface distances on the top, back (occipital lobe), and medial wall of the cortical surfaces. Combining with the manual QC results, the large quantitative differences on the top and back of the cortical surfaces imply that our DL-based pipeline can reduce the systematical errors introduced by original dHCP pipeline in these regions observed in Figure 9. Besides, the large surface distance on the medial wall, which is out of the ROI in cortical surface analysis [23, 26, 50], is expected because original dHCP pipeline extracts cortical surfaces from segmentations that contain the clear boundary of the medial wall [50], whereas our DL-based pipeline extracts cortical surface end-to-end from brain MRI scans.
@@ 4.1.3 Brain Development
We provide qualitative and quantitative evaluation to validate that the cortical surfaces extracted by our dHCP DL pipeline can capture the brain development of neonates. We visualize the cortical surfaces as well as corresponding cortical features of different neonatal subjects scanned at the PMA of 29,34,39,44 weeks in Figure 13. The cortical folding is observed in Figure 13 with the increasing of the PMA. The cortical thickness and sulcal depth are increasing as well with brain growth.
For both original and DL-based dHCP pipeline, we measure and plot cortical morphological features including cortical volume, surface area, cortical thickness and sulcal depth across all 265 test neonatal subjects in Figure 14. The cortical volume is equal to the volumes of pial surfaces minus those of the white matter surfaces for both brain hemispheres. The surface area refers to the mesh area of the pial surfaces. The cortical thickness and sulcal depth for each subject is measured by averaging the cortical thickness and the absolute value of the sulcal depth over all vertices on the surface meshes. Figure 14 shows the positive trends of the morphological features with increasing PMA, which reflect the growth and folding of the neonatal cerebral cortex. Moreover, we compute the Pearson correlation coefficients (PCC) $r$ to measure the correlation between the morphological features and the PMA. The PCCs are provided in Figure 14, which demonstrate that all cortical morphological features have strong positive correlation with the age of the neonates. This verifies that the cortical surfaces extracted by our DL-based pipeline are able to capture the brain development of neonates as the original pipeline.
Compared to the original dHCP pipeline, the cortical volume and surface area measured by our DL-based pipeline manifest similar trends. For cortical thickness, our DL-based pipeline presents stronger correlation and higher PCC value than the original pipeline. We notice that in Figure 14, the absolute value of sulcal depth estimated by the original pipeline is much higher than ours. As mentioned in Section 2.3 and 2.6, the cortical surfaces extracted by original dHCP pipeline have adapted number of vertices, i.e., less vertices for younger subjects. As a consequence, the cortical surfaces of younger infants are over-inflated during the inflation process, thereby causing overestimation of sulcal depth. In contrast, our DL-based pipeline creates cortical surfaces with fixed number of vertices, ensuring age-consistent measurement of the sulcal depth across the PMA.
@@
Table 4: The number of self-intersecting faces (SIFs) on cortical surfaces.